the longer side of a rectangle is double the length of the shorter side. If the length of diagonal is 10cm, what are the dimensions of the rectangle?
Answer by richwmiller(17219)

x^2+(2x)^2=100
x^2+4x^2=100
5x^2=100
x^2=20
x=2sqrt(5)
2x=4sqrt(5)
check
(2sqrt(5))^2+(4sqrt(5))^2=100
4*5+16*5
20+80=100
ok